OpenStudy (anonymous):

find all zeros of the quadratic equation by factoring f(x)=2Xsquared-7X-4

OpenStudy (anonymous):

2x^2 -7x-4 = 0 2x^2 - 8x+x - 4 = 0 2x(x-4)+1(x-4)= 0 (2x+1)(x-4)=0 2x+1 =0 or x-4 = 0 x = -1/2 , x = 4 these are the required zeros

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@rizwan_uet how do i find the zeros of this quadratic equation by completing the square? \[x^{2}+10x+29\]